### v1

@@ 5 Conclusions
This note has extended the previous work of Apers:2022cyl on the relationship between kination and Kasner solutions in type IIB compactifications. Here we have provided much more detail on the kination-Kasner relationship. In the first part of this paper, we extended the earlier analysis of Apers:2022cyl to multiple moduli and other corners of the M-theory duality web, such as type I, heterotic, IIA and also M-theory.
@@ Appendix A Appendix
¯$\begin{split}&\bar{\nabla}_{\mu}H_{\nu\rho}=\nabla_{\mu}h_{\nu\rho}-g_{\nu\rho}\nabla_{\mu}\phi,\\ &\bar{\nabla}_{\mu}H_{\nu i}=\zeta\nabla_{\mu}A_{\nu i}-\tfrac{1}{3}g_{\mu 0}A_{\nu i}\\ &\bar{\nabla}_{\mu}H_{ij}=2\zeta\nabla_{\mu}\phi_{ij},\\ &\bar{\nabla}_{i}H_{\mu\nu}=\partial_{i}h_{\mu\nu}-g_{\mu\nu}\partial_{i}\phi+\tfrac{1}{3}g_{\mu 0}A_{\nu i}+\tfrac{1}{3}g_{\nu 0}A_{\mu i},\\ &\bar{\nabla}_{i}H_{j\mu}=\zeta\partial_{i}A_{\mu j}+\tfrac{1}{3}g_{\mu 0}[\phi\zeta^{-1}g_{ij}+2\phi_{ij}],\\ &\bar{\nabla}_{i}H_{jk}=2\zeta\partial_{i}\phi_{jk}.\end{split}$ (75)
@@
¯$\begin{split}&\bar{\nabla}^{\mu}H^{\nu\rho}=\nabla^{\mu}h^{\nu\rho}-g^{\nu\rho}\nabla^{\mu}\phi,\\ &\bar{\nabla}^{\mu}H^{\nu i}=\zeta\nabla^{\mu}A^{\nu i}-\delta^{\mu}_{0}A^{\nu i}=\zeta g^{ij}\nabla^{\mu}A^{\nu}_{j}-\tfrac{1}{3}\delta^{\mu}_{0}A^{\nu i}\\ &\bar{\nabla}^{\mu}H^{ij}=2\zeta g^{ik}g^{jm}\nabla^{\mu}\phi_{km},\\ &\bar{\nabla}^{i}H^{\mu\nu}=\partial^{i}h^{\mu\nu}-g^{\mu\nu}\partial^{i}\phi+\tfrac{1}{3}\delta^{\mu}_{0}A^{\nu i}+\tfrac{1}{3}\delta^{\nu}_{0}A^{\mu i},\\ &\bar{\nabla}^{i}H^{j\mu}=\zeta\partial^{i}A^{\mu j}+\tfrac{1}{3}\delta^{\mu}_{0}[\phi\zeta^{-1}g^{ij}+2\phi^{ij}],\\ &\bar{\nabla}^{i}H^{jk}=2\zeta\partial^{i}\phi^{jk}.\end{split}$ (76)
@@
$-\tfrac{1}{4}\bar{\nabla}_{M}H\bar{\nabla}^{M}H=-\tfrac{1}{4}\partial_{\mu}h\partial^{\mu}h+\partial_{\mu}h\partial^{\mu}\phi-\partial_{\mu}\phi\partial^{\mu}\phi-\tfrac{1}{4}\partial_{i}h\partial^{i}h+\partial_{i}h\partial^{i}\phi-\partial_{i}\phi\partial^{i}\phi.$ (77)
It is also useful to evaluate
$\begin{split}\tfrac{1}{2}\bar{\nabla}_{M}H^{MN}\bar{\nabla}_{N}H&=\tfrac{1}{2}\nabla_{\mu}h^{\mu\alpha}\nabla_{\alpha}h-\nabla_{\mu}h^{\mu\alpha}\nabla_{\alpha}\phi-\tfrac{1}{2}\nabla_{\mu}h\nabla^{\mu}\phi+\nabla_{\mu}\phi\nabla^{\mu}\phi\\ &+\tfrac{4}{3}\zeta^{-1}\phi\nabla_{0}h-\tfrac{8}{3}\zeta^{-1}\phi\nabla_{0}\phi\\ &+\tfrac{1}{2}\zeta\nabla_{\mu}A^{\mu}_{i}\partial^{i}h-\zeta\nabla_{\mu}A^{\mu}_{i}\partial^{i}\phi+\tfrac{1}{2}\zeta\partial_{i}A^{\mu i}\nabla_{\mu}h-\zeta\partial_{i}A^{\mu i}\nabla_{\mu}\phi\\ &+\zeta\partial_{i}\phi^{ij}\partial_{j}h-2\zeta\partial_{i}\phi^{ij}\partial_{j}\phi,\end{split}$ (78)
@@
$\begin{split}-\tfrac{1}{2}\bar{\nabla}_{M}H_{AB}\bar{\nabla}^{A}H^{MB}&=-\tfrac{1}{2}\nabla_{\mu}h_{\nu\rho}\nabla^{\nu}h^{\mu\rho}+\nabla_{\mu}h^{\mu\nu}\nabla_{\nu}\phi-\tfrac{1}{2}\nabla_{\mu}\phi\nabla^{\mu}\phi-\tfrac{1}{2}\zeta^{2}g^{ij}\nabla_{\mu}A_{\nu i}\nabla^{\nu}A^{\mu}_{j}\\ &-\tfrac{1}{9}\zeta^{-1}A_{\nu i}A^{\nu i}+\tfrac{1}{9}\zeta^{-1}[5\zeta^{-2}\phi^{2}+2\phi_{ij}\phi^{ij}]\\ &-\tfrac{1}{3}\zeta A^{\nu i}\nabla_{0}A_{\nu i}-\tfrac{2}{3}\zeta^{-1}\phi\nabla_{0}\phi-\tfrac{4}{3}\zeta\phi^{ij}\nabla_{0}\phi_{ij}\\ &-2\zeta^{2}\partial_{i}\phi_{jk}\partial^{j}\phi^{ik}-\tfrac{1}{2}\zeta^{2}\partial_{i}A_{\mu j}\partial^{j}A^{\mu i}\\ &-\zeta\nabla_{\mu}A_{\nu i}\partial^{i}h^{\mu\nu}+\zeta\partial^{i}\phi\nabla_{\mu}A^{\mu}_{i}-2\zeta^{2}\nabla_{\mu}\phi_{ij}\partial^{i}A^{\mu j},\end{split}$ (79)
@@
$\begin{split}\tfrac{1}{4}\bar{\nabla}_{M}H_{AB}\bar{\nabla}^{M}H^{AB}&=\tfrac{1}{4}\nabla_{\mu}h_{\nu\rho}\nabla^{\mu}h^{\nu\rho}-\tfrac{1}{2}\nabla_{\mu}h\nabla^{\mu}\phi\\ &+\nabla_{\mu}\phi\nabla^{\mu}\phi+\tfrac{1}{2}\zeta^{2}g^{ij}\nabla_{\mu}A_{\nu i}\nabla^{\mu}A^{\nu}_{j}+\zeta^{2}g^{ik}g^{jp}\nabla_{\mu}\phi_{ij}\nabla^{\mu}\phi_{kp}\\ &-\tfrac{1}{9}\zeta^{-1}A_{\nu i}A^{\nu i}-\tfrac{1}{9}\zeta^{-1}[5\phi^{2}\zeta^{-2}+2\phi_{ij}\phi^{ij}]-\tfrac{1}{3}\zeta A^{\nu i}\nabla_{0}A_{\nu i}\\ &+\tfrac{1}{4}\partial_{i}h_{\mu\nu}\partial^{i}h^{\mu\nu}-\tfrac{1}{2}\partial_{i}h\partial^{i}\phi+\partial_{i}\phi\partial^{i}\phi+\tfrac{1}{2}\zeta^{2}\partial_{i}A_{\mu j}\partial^{i}A^{\mu j}+\zeta^{2}\partial_{i}\phi_{jk}\partial^{i}\phi^{jk}.\end{split}$ (80)

### v2

@@ 5 Conclusions
This note has extended the previous work of Apers et al. [2023] on the relationship between kination and Kasner solutions in type IIB compactifications. Here we have provided much more detail on the kination-Kasner relationship. In the first part of this paper, we extended the earlier analysis of Apers et al. [2023] to multiple moduli and other corners of the M-theory duality web, such as type I, heterotic, IIA and also M-theory.
@@ Appendix A Appendix
$\begin{split}&\bar{\nabla}_{\mu}H_{\nu\rho}=\nabla_{\mu}h_{\nu\rho}-g_{\nu\rho}\nabla_{\mu}\phi,\\ &\bar{\nabla}_{\mu}H_{\nu i}=\zeta\nabla_{\mu}A_{\nu i}-\tfrac{1}{3}g_{\mu 0}A_{\nu i}\\ &\bar{\nabla}_{\mu}H_{ij}=2\zeta\nabla_{\mu}\phi_{ij},\\ &\bar{\nabla}_{i}H_{\mu\nu}=\partial_{i}h_{\mu\nu}-g_{\mu\nu}\partial_{i}\phi+\tfrac{1}{3}g_{\mu 0}A_{\nu i}+\tfrac{1}{3}g_{\nu 0}A_{\mu i},\\ &\bar{\nabla}_{i}H_{\mu j}=\zeta\partial_{i}A_{\mu j}+\tfrac{1}{3}g_{\mu 0}[2\phi_{ij}+\zeta^{-1}g_{ij}\phi]-\tfrac{1}{3}\zeta^{-1}g_{ij}h_{0\mu},\\ &\bar{\nabla}_{i}H_{jk}=2\zeta\partial_{i}\phi_{jk}-\tfrac{1}{3}g_{ij}A_{0k}-\tfrac{1}{3}g_{ik}A_{0j}.\end{split}$ (77)
@@
[…]
where Latin indices were raised with $g^{ij}$ while Greek ones with $g^{\mu\nu}$. We have to be careful not to commute $g^{ij}$ past $\nabla_{\mu}$ covariant derivatives since only the non-compact metric commutes with these.
@@
$\tfrac{1}{4}\bar{\nabla}_{M}H\bar{\nabla}^{M}H=-\tfrac{1}{4}\partial_{\mu}h\partial^{\mu}h+\partial_{\mu}h\partial^{\mu}\phi-\partial_{\mu}\phi\partial^{\mu}\phi-\tfrac{1}{4}\partial_{i}h\partial^{i}h+\partial_{i}h\partial^{i}\phi-\partial_{i}\phi\partial^{i}\phi.$ (79)
It is also useful to evaluate, up to integration by parts, that
$\begin{split}\tfrac{1}{2}\bar{\nabla}_{M}H^{MN}\bar{\nabla}_{N}H&=\tfrac{1}{2}\nabla_{\mu}h^{\mu\nu}\nabla_{\nu}h-\nabla_{\mu}h^{\mu\nu}\nabla_{\nu}\phi-\tfrac{1}{2}\nabla_{\mu}h\nabla^{\mu}\phi+\nabla_{\mu}\phi\nabla^{\mu}\phi\\ &+\tfrac{4}{3}\zeta^{-1}\phi\nabla_{0}h-\tfrac{8}{3}\zeta^{-1}\phi\nabla_{0}\phi-\zeta^{-1}h_{0}^{\mu}\nabla_{\mu}h+2\zeta^{-1}h_{0}^{\mu}\nabla_{\mu}\phi\\ &-\tfrac{4}{3}A_{0}^{i}\partial_{i}h+\tfrac{8}{3}A_{0}^{i}\partial_{i}\phi\\ &+\tfrac{1}{2}\zeta\nabla_{\mu}A^{\mu}_{i}\partial^{i}h-\zeta\nabla_{\mu}A^{\mu}_{i}\partial^{i}\phi+\tfrac{1}{2}\zeta\partial_{i}A^{\mu i}\nabla_{\mu}h-\zeta\partial_{i}A^{\mu i}\nabla_{\mu}\phi\\ &-\zeta\partial_{i}\phi^{ij}\partial_{j}h+2\zeta\partial_{i}\phi^{ij}\partial_{j}\phi,\end{split}$ (80)
@@
$\begin{split}-\tfrac{1}{2}\bar{\nabla}_{M}H_{AB}\bar{\nabla}^{A}H^{MB}&=-\tfrac{1}{2}\nabla_{\mu}h_{\nu\rho}\nabla^{\nu}h^{\mu\rho}+\nabla_{\mu}h^{\mu\nu}\nabla_{\nu}\phi-\tfrac{1}{2}\nabla_{\mu}\phi\nabla^{\mu}\phi-\tfrac{1}{2}\zeta^{2}g^{ij}\nabla_{\mu}A_{\nu i}\nabla^{\nu}A^{\mu}_{j}\\ &-\tfrac{1}{9}\zeta^{-1}A_{\mu i}A^{\mu i}-\tfrac{4}{9}A_{0i}A_{0}{}^{i}+\tfrac{1}{9}\zeta^{-1}[5\zeta^{-2}\phi^{2}+2\phi_{ij}\phi^{ij}]+\tfrac{8}{9}\zeta^{-2}h_{00}\phi-\tfrac{1}{3}\zeta^{-2}h_{0\mu}h_{0}{}^{\mu}\\ &-\tfrac{1}{3}\zeta A^{\mu i}\nabla_{0}A_{\mu i}-\tfrac{4}{3}\zeta\phi^{ij}\nabla_{0}\phi_{ij}-\tfrac{2}{3}g^{ij}\phi\nabla_{0}\phi_{ij}+\tfrac{2}{3}h_{0}^{\mu}g^{ij}\nabla_{\mu}\phi_{ij}\\ &-\tfrac{2}{3}\partial^{i}\phi A_{0i}+\tfrac{4}{3}\zeta\partial_{i}\phi^{ij}A_{0j}\\ &-2\zeta^{2}\partial_{i}\phi_{jk}\partial^{j}\phi^{ik}-\tfrac{1}{2}\zeta^{2}\partial_{i}A_{\mu j}\partial^{j}A^{\mu i}\\ &-\zeta\nabla_{\mu}A_{\nu i}\partial^{i}h^{\mu\nu}+\zeta\partial^{i}\phi\nabla_{\mu}A^{\mu}_{i}-2\zeta^{2}\nabla_{\mu}\phi_{ij}\partial^{i}A^{\mu j},\end{split}$ (81)
@@
$\begin{split}\tfrac{1}{4}\bar{\nabla}_{M}H_{AB}\bar{\nabla}^{M}H^{AB}&=\tfrac{1}{4}\nabla_{\mu}h_{\nu\rho}\nabla^{\mu}h^{\nu\rho}-\tfrac{1}{2}\nabla_{\mu}h\nabla^{\mu}\phi+\nabla_{\mu}\phi\nabla^{\mu}\phi\\ &+\tfrac{1}{2}\zeta^{2}g^{ij}\nabla_{\mu}A_{\nu i}\nabla^{\mu}A^{\nu}_{j}+\zeta^{2}g^{ik}g^{jp}\nabla_{\mu}\phi_{ij}\nabla^{\mu}\phi_{kp}\\ &-\tfrac{1}{9}\zeta^{-1}A_{\mu i}A^{\mu i}+\tfrac{4}{9}A_{0i}A_{0}{}^{i}-\tfrac{1}{9}\zeta^{-1}[5\zeta^{-2}\phi^{2}+2\phi_{ij}\phi^{ij}]-\tfrac{8}{9}\zeta^{-2}h_{00}\phi+\tfrac{1}{3}\zeta^{-2}h_{0\mu}h^{\mu}_{0}\\ &-\tfrac{1}{3}\zeta A^{\mu i}\nabla_{0}A_{\mu i}-\tfrac{2}{3}A_{0i}\partial^{i}\phi+\tfrac{4}{3}\zeta\phi^{ij}\partial_{i}A_{0j}-\tfrac{2}{3}h_{0}^{\mu}\partial^{i}A_{\mu i}\\ &+\tfrac{1}{4}\partial_{i}h_{\mu\nu}\partial^{i}h^{\mu\nu}-\tfrac{1}{2}\partial_{i}h\partial^{i}\phi+\partial_{i}\phi\partial^{i}\phi+\tfrac{1}{2}\zeta^{2}\partial_{i}A_{\mu j}\partial^{i}A^{\mu j}+\zeta^{2}\partial_{i}\phi_{jk}\partial^{i}\phi^{jk}.\end{split}$ (82)
Here we can freely integrate the compact coordinate by parts since the compact metric $\zeta\delta_{ij}$ is flat, and so all of its Christoffel symbols vanish, and so $\nabla_{i}=\partial_{i}$.
Expansion of the Kination Action
The kination action can be written as
$S=\frac{1}{2\kappa_{4}^{2}}\int d^{4}x\sqrt{-g^{\prime}}\bigg[R^{\prime}-\frac{1}{2}g^{\prime\mu\nu}\partial_{\mu}\Phi\partial_{\nu}\Phi\bigg],$ (83)
where the scalar field can be canonically normalized through $\Phi\rightarrow\sqrt{2}M_{P}^{-1}\Phi$. To match its perturbations to those acquired from the compactification of the Kasner solution in Eq. (49), we need to expand this solution to second order in the perturbations $g_{\mu\nu}^{\prime}=g_{\mu\nu}+h_{\mu\nu}$ and $\Phi=\bar{\Phi}+\varphi$. Here the kination metric is given by
$g_{\mu\nu}=\tfrac{4\eta}{3}\eta_{\mu\nu},$ (84)
for which
$R_{\mu\nu}=\tfrac{3}{2\eta^{2}}\delta_{\mu 0}\delta_{\nu 0},\ \ \ \ \ \ R=-\tfrac{9}{8\eta^{3}}.$ (85)
Meanwhile,
$\bar{\Phi}=\bar{\Phi}_{0}+\sqrt{3}\ln(\eta/\eta_{0}).$ (86)
It is useful to introduce the vector field $B_{\mu}=\partial_{\mu}\bar{\Phi}=\tfrac{\sqrt{3}}{\eta}\delta_{\mu 0}$, which satisfies $R_{\mu\nu}=B_{\mu}B_{\nu}/2$ and $R=B^{\mu}B_{\mu}/2$.
To second order in metric perturbations the determinant is expanded out as Ortín [2015]
$\sqrt{-g^{\prime}}=\sqrt{-g}(1+\tfrac{1}{2}h+\tfrac{1}{8}h^{2}-\tfrac{1}{4}h_{\mu\nu}h^{\mu\nu}+\cdots),$ (87)
and Ricci scalar as $R^{\prime}=R^{(0)}+R^{(1)}+R^{(2)}$, where
$\begin{split}R^{(0)}&=R,\\ R^{(1)}&=-h^{\mu\nu}R_{\mu\nu}+\nabla_{\mu}\nabla_{\nu}h^{\mu\nu}-\nabla_{\mu}\nabla^{\mu}h,\\ R^{(2)}&=2h^{\mu}{}_{\rho}h^{\rho\nu}R_{\mu\nu}-h^{\mu\nu}h^{\rho\lambda}R_{\mu\rho\nu\lambda}+h^{\mu\nu}\nabla_{\mu}\nabla_{\nu}h-\tfrac{1}{4}\nabla_{\mu}h\nabla^{\mu}h-\nabla_{\mu}h^{\mu\nu}\nabla_{\rho}h_{\nu}{}^{\rho}\\ &+\nabla_{\mu}h^{\mu\nu}\nabla_{\nu}h-2h^{\mu\nu}\nabla_{\rho}\nabla_{\mu}h_{\nu}{}^{\rho}+h_{\mu\nu}\nabla_{\rho}\nabla^{\rho}h^{\mu\nu}-\tfrac{1}{2}\nabla_{\mu}h_{\nu\rho}\nabla^{\rho}h^{\mu\nu}+\tfrac{3}{4}\nabla_{\rho}h_{\mu\nu}\nabla^{\rho}h^{\mu\nu}.\end{split}$ (88)
The Einstein–Hilbert Lagrangian can then be expressed as
$\frac{1}{2\kappa_{4}^{2}}\frac{\sqrt{-g^{\prime}}}{\sqrt{-g}}R^{\prime}=R^{(0)}+[R^{(1)}+\tfrac{1}{2}hR^{(0)}]+[(\tfrac{1}{8}h^{2}-\tfrac{1}{4}h_{\mu\nu}h^{\mu\nu})R^{(0)}+\tfrac{1}{2}hR^{(1)}+R^{(2)}].$ (89)
The leading order term is the background Einstein–Hilbert Lagrangian
$\mathcal{L}_{h}^{(0)}=\frac{\sqrt{-g}}{2\kappa_{4}^{2}}R,$ (90)
while, by dropping total derivatives, the linear term can be written as
$\mathcal{L}_{h}^{(1)}=\frac{\sqrt{-g}}{2\kappa_{4}^{2}}\Big[\tfrac{1}{2}hR-h^{\mu\nu}R_{\mu\nu}\Big],$ (91)
Using
$\nabla_{\mu},\nabla_{\nu}]h^{\alpha}{}_{\beta}=R^{\alpha}{}_{\rho\mu\nu}h^{\rho}{}_{\beta}-R^{\rho}{}_{\beta\mu\nu}h^{\alpha}{}_{\rho}$, (92)
the $\sqrt{-g}R^{(2)}$ part of the second order term can be written up to total derivatives as
$\sqrt{-g}R^{(2)}=\sqrt{-g}[h^{\mu}{}_{\rho}h^{\rho\nu}R_{\mu\nu}-\tfrac{1}{4}\nabla_{\mu}h\nabla^{\mu}h-\tfrac{1}{4}\nabla_{\rho}h_{\mu\nu}\nabla^{\rho}h^{\mu\nu}+\tfrac{1}{2}\nabla_{\rho}h^{\mu\nu}\nabla_{\mu}h_{\nu}{}^{\rho}$]. (93)
Combining this with the other second-order components gives
$\begin{split}\mathcal{L}_{h}^{(2)}&=\frac{\sqrt{-g}}{2\kappa_{4}^{2}}\bigg[\tfrac{1}{4}\nabla_{\mu}h\nabla^{\mu}h-\tfrac{1}{4}\nabla_{\rho}h_{\mu\nu}\nabla^{\rho}h^{\mu\nu}+\tfrac{1}{2}\nabla_{\rho}h^{\mu\nu}\nabla_{\mu}h_{\nu}{}^{\rho}-\tfrac{1}{2}\nabla_{\mu}h\nabla_{\nu}h^{\mu\nu}\\ &+\tfrac{1}{2}h^{\mu}{}_{\rho}h^{\rho\nu}R_{\mu\nu}+\tfrac{1}{2}h^{\mu\nu}h^{\rho\lambda}R_{\mu\rho\nu\lambda}-\tfrac{1}{2}hh^{\mu\nu}R_{\mu\nu}+\tfrac{1}{8}h^{2}R-\tfrac{1}{4}h_{\mu\nu}h^{\mu\nu}R\bigg].\end{split}$ (94)
The kination field Lagrangian can be expanded in perturbations and written as $\mathcal{L}_{\varphi}=\mathcal{L}_{\varphi}^{(0)}+\mathcal{L}_{\varphi}^{(1)}+\mathcal{L}_{\varphi}^{(2)}$ where
$\begin{split}\mathcal{L}_{\varphi}^{(0)}&=-\frac{\sqrt{-g}}{2\kappa_{4}^{2}}R,\\ \mathcal{L}_{\varphi}^{(1)}&=-\frac{\sqrt{-g}}{2\kappa_{4}^{2}}\bigg[X^{\mu}\partial_{\mu}\varphi-h^{\mu\nu}R_{\mu\nu}+\tfrac{1}{2}hR\bigg],\\ \mathcal{L}^{(2)}_{\varphi}&=-\frac{\sqrt{-g}}{2\kappa_{4}^{2}}\bigg[\tfrac{1}{8}Rh^{2}-\tfrac{1}{4}Rh_{\mu\nu}h^{\mu\nu}+h^{\mu}{}_{\rho}h^{\rho\nu}R_{\mu\nu}-\tfrac{1}{2}hh^{\mu\nu}R_{\mu\nu}\\ &\ \ \ \ \ \ \ \ \ \ \ \ \ \ +\tfrac{1}{2}\partial^{\mu}\varphi\partial_{\mu}\varphi-h^{\mu\nu}B_{\mu}\partial_{\nu}\varphi+\tfrac{1}{2}hB^{\mu}\partial_{\mu}\varphi\bigg].\end{split}$ (95)
The total Lagrangian at each order in the perturbations is given by combining the contributions from the Einstein–Hilbert term and the scalar kinetic term $\mathcal{L}^{(i)}=\mathcal{L}^{(i)}_{h}+\mathcal{L}^{(i)}_{\varphi}$. The leading order term $\mathcal{L}^{(0)}$ vanishes since the background kination solution is a solution to the Einstein equation. Meanwhile, most linear Lagrangian terms cancel with the remaining one being a total derivative
$\mathcal{L}^{(1)}=-\frac{\sqrt{-g}}{2\kappa_{4}^{2}}B^{\mu}\partial_{\mu}\varphi=-\frac{\sqrt{-g}}{2\kappa_{4}^{2}}\nabla_{\mu}(B^{\mu}\varphi).$ (96)
The only non-vanishing terms are found in the bilinear Lagrangian
@@
$\begin{split}\mathcal{L}^{(2)}&=\frac{\sqrt{-g}}{2\kappa_{4}^{2}}\bigg[\tfrac{1}{4}\nabla_{\mu}h\nabla^{\mu}h-\tfrac{1}{4}\nabla_{\rho}h_{\mu\nu}\nabla^{\rho}h^{\mu\nu}+\tfrac{1}{2}\nabla_{\mu}h^{\mu\nu}\nabla_{\rho}h_{\nu}{}^{\rho}-\tfrac{1}{2}\nabla_{\mu}h\nabla_{\nu}h^{\mu\nu}\\ &-\tfrac{1}{2}\partial^{\mu}\varphi\partial_{\mu}\varphi+h^{\mu\nu}B_{\mu}\partial_{\nu}\varphi-\tfrac{1}{2}hB^{\mu}\partial_{\mu}\varphi\bigg].\end{split}$ (97)
This matches the Lagrangian for the metric perturbations and the dilaton acquired from the compactification of perturbations of the Kasner solution in Eq. (49).